OpenStudy (anonymous):

how would i solve this ? 6 − x = 15

OpenStudy (zzr0ck3r):

start by subtracting 6 from woth sides and see what you get

OpenStudy (anonymous):

what you need to do is isolate x, so that it is positive and by itself this can be done by subtracting 6 from both sides of the equation \[6-x-6=15-6\] \[-x=9\] because x needs to be positive, this can be done by multiplying it by (-1) however, whenever you do anything to one side of the equation, the same must be done for the other side \[(-1)*(-x)=(-1)*(9)\] \[x=-9\] for any future problems, always remember that if you ever add, subtract, multiply, divide. you must always do it to both sides

OpenStudy (anonymous):

6-6 is0 and 15-6=9

OpenStudy (anonymous):

so would it be x=9-15 or some thing ?